# Forest Proximities for Time Series

### v2

@@ V Discussion of Limitations
Another possible limitation in our method is the formula we have used to define dissimilarities from $p_{GAP}$, given in Equation 5. We have advocated for the use of the exponent value of $2$ on the basis that the proximities are typically small. However, it is possible that using a different formula to define dissimilarity from $p_{GAP}$ proximities will have a positive or negative impact on subsequent tasks, including obtaining time series embeddings and outlier detection.
@@
There may also be limitations imposed by the proximity forest classifier itself. The number of proximity trees chosen is a hyperparameter and is likely to have a large impact on the quality of the $p_{GAP}$ proximities. In particular, random forests are commonly trained with several hundred trees, whereas a tree count of less than $100$ trees is not uncommon for a proximity forest. When fewer trees are chosen, the likelihood of points being in-bag for every tree increases, affecting the proximities $p_{GAP}$.
VI Conclusion
@@
We have introduced forest proximities for proximity forests, specifically $p_{GAP}$ proximities previously defined for random forests [2], which we term PF-GAP. Using the forest proximities, we have constructed time series vector embeddings. We show that they seem to produce better inter-class separation than unsupervised DTW-induced embeddings using three examples. We have also introduced intraclass outlier scores for time series based on forest proximities. With the GunPoint and ArrowHead datasets [20], we have visually shown that the time series outlier scores may correspond with outliers in the embedded space. We have also used the forest proximities along with LOF to detect outliers, showing a possible connection between points which are misclassified by a proximity forest classifier and points which are considered outliers using the forest proximities. Future work includes further study of the use of PF-GAP in time series outlier detection–that is, in the detection of within-class anomalous time series.
@@
Future work also includes the exploration of additional applications of forest proximities. With random forests, forest proximities have been used directly in classification without the need to predict using the forest structure [2]. This and other potential applications of forest proximities for proximity forests we leave as a future endeavor. Additionally, a “proximity forest 2.0” algorithm has recently been introduced [21], promising improved computational speed and classification accuracy over the original proximity forest algorithm. Thus, it is natural to consider the future work of calculating forest proximities for proximity forest 2.0.

### v3

@@ 5 Discussion of Limitations
One possible limitation in our method is the formula we have used to define dissimilarities from $p_{GAP}$, given in Equation 5. We have advocated for the use of the exponent value of $2$ on the basis that the proximities are typically small. However, it is possible that using a different formula to define dissimilarity from $p_{GAP}$ proximities will have a positive or negative impact on subsequent tasks, including obtaining time series embeddings and outlier detection.
@@
There may also be limitations imposed by the proximity forest classifier itself. The number of proximity trees chosen is a hyperparameter and is likely to have a large impact on the quality of the $p_{GAP}$ proximities. In particular, random forests are commonly trained with several hundred trees, whereas a tree count of less than $100$ trees is not uncommon for a proximity forest. When fewer trees are chosen, the likelihood of points being in-bag for every tree increases, affecting the proximities defined by $p_{GAP}$.
6 Conclusion
@@
We have introduced forest proximities for the PF model, specifically $p_{GAP}$ proximities previously defined for random forests [2], which we term PF-GAP. Using the forest proximities, we have demonstrated how to construct time series vector embeddings, which may produce better inter-class separation than unsupervised time series distance-induced embeddings. We have also introduced intraclass outlier scores for time series based on forest proximities. We have used the forest proximities along with LOF to detect outliers, showing a strong connection between points which are misclassified by a proximity forest classifier and points which are considered outliers using the forest proximities. Future work includes further study of the use of PF-GAP in time series outlier detection–that is, in the detection of within-class anomalous time series.
@@
Future work also includes the exploration of additional applications of forest proximities. With random forests, forest proximities have been used directly in classification without the need to predict using the forest structure [2]. This and other potential applications of forest proximities for proximity forests we leave as a future endeavor. Additionally, a “proximity forest 2.0” algorithm has recently been introduced [22], promising improved computational speed and classification accuracy over the original proximity forest algorithm. Thus, it is natural to consider the future work of calculating forest proximities for proximity forest 2.0, as well as for other forest-based time series classifiers.